solve 5x+x(x-1)=0 and find the x-intercepts of f(x)= 5x+x(x-1)
Answer by stanbon(57307)

solve 5x+x(x-1)=0 and find the x-intercepts of f(x)= 5x+x(x-1)
--------------------
5x+x(x-1)=0
---
5x + x^2 -x = 0
x^2+4x = 0
Factor:
x(x+4) = 0
x = 0 or x = -4
---
They ARE the x-intercepts.
Cheers,
Stan H.